### v1

# Wildfire Risk Prediction: A Review

Abstract
Wildfires have significant impacts on global vegetation, wildlife, and humans. They destroy plant communities and wildlife habitats and contribute to increased emissions of carbon dioxide, nitrogen oxides, methane, and other pollutants. The prediction of wildfires relies on various independent variables combined with regression or machine learning methods. In this technical review, we describe the options for independent variables, data processing techniques, models, independent variables collinearity and importance estimation methods, and model performance evaluation metrics. First, we divide the independent variables into 4 aspects, including climate and meteorology conditions, socio-economical factors, terrain and hydrological features, and wildfire historical records. Second, preprocessing methods are described for different magnitudes, different spatial-temporal resolutions, and different formats of data. Third, the collinearity and importance evaluation methods of independent variables are also considered. Fourth, we discuss the application of statistical models, traditional machine learning models, and deep learning models in wildfire risk prediction. In this subsection, compared with other reviews, this manuscript particularly discusses the evaluation metrics and recent advancements in deep learning methods. Lastly, addressing the limitations of current research, this paper emphasizes the need for more effective deep learning time series forecasting algorithms, the utilization of three-dimensional data including ground and trunk fuel, extraction of more accurate historical fire point data, and improved model evaluation metrics.
keywords:
wildfire prediction , remote sensing , machine learning , review
[…]
1 Introduction
Wildfires can occur in various environments, such as forests, grasslands, and peatlands, exerting significant impacts on global vegetation, wildlife, and human health, as well as air quality and carbon balance. They destroy plant communities and wildlife habitats, leading to increased emissions of greenhouse gases such as carbon dioxide, nitrogen oxides, methane ozone, as well as fine particulate matter, and mercury (Pelletier et al., 2023; Hua and Shao, 2017; McNorton and Di Giuseppe, 2024; Yang et al., 2022). For instance, in 2019 alone, widespread forest fires in Australia burned approximately 18.6 million hectares of land, destroyed over 5,900 structures, and resulted in the deaths of at least 34 people and one billion animals (Naderpour et al., 2020). In the coming decades, climate change may shorten fire return intervals and increase the burned area in vulnerable regions (Parks and Abatzoglou, 2020; Holdrege et al., 2024). Therefore, accurate wildfire risk prediction is of utmost importance.
Since the 1980s, wildfire prediction has gained increasing attention from governments and academia, leading to the development of wildfire prediction models tailored to different geographic regions. Common wildfire risk prediction models encompass various aspects, including the determination of study area and temporal scales, data selection and processing, method selection, identification of prediction objectives, and model accuracy evaluation. However, there is still a lack of systematic review regarding these key components of research. Thus, this paper systematically reviews and organizes the aforementioned aspects, aiming to support the selection of appropriate data, data processing procedures, and wildfire risk prediction methods for different ecological zones and land cover types. Furthermore, this manuscript also provides an outlook on potential directions for future research in wildfire risk prediction.
Firstly, wildfire prediction models require independent variables and dependent variables, i.e., the input data required by the model training and ground truth data used for training and accuracy validation. The data generally encompass five aspects: fuel conditions, climate and meteorology conditions, socio-economic factors, topographic and hydrological features, socio-economic factors, and wildfire historical records. Due to the varying magnitudes and spatiotemporal resolutions of different types of data used in research, such as meteorological data, remote sensing indices, and socio-economic data, this manuscript will also introduce some data preprocessing methods commonly employed in studies. Moreover, the evaluation of collinearity and importance among different features is crucial during algorithm design, as including strongly correlated features simultaneously in the model can lead to performance degradation and interfere with feature importance assessment. This manuscript will present several techniques for assessing the collinearity and significance of features.
Another important aspect of wildfire prediction models is the selection of algorithms. Currently, commonly used approaches in wildfire prediction include fire danger rating systems, statistical models, and traditional machine learning models. Furthermore, the rapid development of deep learning has also attracted the attention of researchers in the field of wildfire prediction. Firstly, we briefly describe the commonly utilized wildfire danger rating systems. Then, we introduced commonly used statistical models including Logistic Regression Models (LR), Generalized Linear Models (GLM), Geographically weighted logistic regression models, and others. Thirdly, commonly used machine learning models for wildfire prediction are not included in this manuscript as Jain et al. (2020) have systemically discussed them before. In recent years, rapidly evolving deep learning models have demonstrated superior performance over traditional machine learning models in various tasks, making them important tools for wildfire prediction. However, a comprehensive summary of their applications in wildfire prediction is currently lacking. Therefore, in this study, based on the different deep learning models employed in wildfire prediction, they are categorized as recurrent neural network (RNN) (including long short-term memory (LSTM) and gated recurrent unit (GRU)), convolutional neural network (CNN), graph neural network (GNN), graph convolutional neural network (GCN), and Transformer.
2 Data
As mentioned earlier, this study divides the input variables into five aspects: fuel conditions, meteorology and climate conditions, socio-economic factors, terrain variables and hydrological features, and wildfire historical records.
2.1 Fuel conditions
Fuel condition mainly includes fuel moisture content (FMC), fuel load, fuel type, and fuel continuity. These parameters directly influence the probability, intensity, and spread rate of fires. It is worth noting that fuel type and load provide the material for wildfire combustion and are the main driving factors of wildfire severity (Walker et al., 2020).
2.1.1 Fuel moisture content
Fuel moisture content refers to the ratio of water content to fresh leaf weight or dry leaf weight (Mbow et al., 2004; Chuvieco et al., 1999), and it regulates the combustibility of fuel (Yebra et al., 2013). Beverly and Wotton (2007) conducted experimental research in Canada and found that the moisture of fine fuels is a primary driver of sustained flaming ignition. Similar results were observed by Leonard (2009) in experiments conducted in Tasmanian grasslands.
Fuel moisture content can be divided into two categories: live fuel moisture content (LFMC) and dead fuel moisture content (DFMC). LFMC is controlled by various eco-physiological characteristics and environmental factors influenced by transpiration, such as soil moisture (Castro et al., 2003). DFMC is unrelated to transpiration and is determined by meteorological factors, including precipitation, solar radiation, fuel diameter, and biochemical composition, which are empirically determined (Viney, 1991; Ager et al., 2010). Current operational fire prediction systems and fire weather indices (FWI) typically rely on estimates of DFMC as a representative of combustibility (Vitolo et al., 2020). Nelson Jr (2000) estimated daily DFMC estimates from 2010 to 2019 at approximately 9 km using the Nelson model, which is a physical model based on the 10-hour fuel moisture content. This model was developed using hourly input of temperature, humidity, radiation, and precipitation, and it predicts fuel moisture content by calculating heat conduction and moisture transport between the wooden cylinder, the surrounding atmosphere, and the exterior. McNorton and Di Giuseppe (2024) designed a new method to estimate DFMC by grouping fuels and simplifying the Nelson model. Additionally, the Cellulose Absorption Index (CAI), Lignin-Cellulose Absorption Index (LCAI), Normalized Difference Tillage Index (NDTI), and Shortwave-Infrared Normalized Difference Residue Index (SINDRI) are commonly used for DFMC estimation (Zacharakis and Tsihrintzis, 2023).
Compared to DFMC, most studies focus more on LFMC because it exhibits significant variations with changes in geography, ecological environment, and vegetation types. In some studies, fuel moisture content is obtained through field measurements and extrapolated to the entire study area through interpolation or remote sensing inversion (Bianchi and Defossé, 2015). However, this approach is time-consuming, labor-intensive, and does not guarantee extrapolation accuracy for large area. In addition, simple physical models are not suitable for heterogeneous vegetation-covered real-world environments. Remote sensing technology, as a low-cost means of large-scale and long-term dynamic Earth observation, can also be used to estimate fuel moisture content in highly heterogeneous real-world scenarios (Quan et al., 2021; Zhu et al., 2021; Verbesselt et al., 2002). Many studies have shown that there is a water absorption band in the near-infrared and shortwave-infrared regions, approximately from 1 µm to 2.5 µm, which is a key parameter for short-term fire hazard estimation (Zacharakis and Tsihrintzis, 2023; Kriedemann and Barrs, 1983). Based on this observation, most studies indirectly estimate vegetation moisture content by observing the spectral reflectance of objects (i.e., spectral features) and constructing specific proxy indices that model the relationship between fuel moisture and vegetation greenness index. Common remote sensing indices used for fuel moisture inversion include the normalized difference vegetation index (NDVI), enhanced vegetation index (EVI), normalized difference infrared index (NDII), water index (WI) (PEÑUELAS et al., 1993), normalized difference water index (NDWI) (Gao, 1996), and shortwave infrared water stress index (SIWSI) (Fensholt and Sandholt, 2003), among others.
Specifically, Malik et al. (2021) predicted the wildfire risk in approximately 63 $km^{2}$ of densely vegetated areas between Monticello and Winters in California. They used Landsat 8 data to construct NDVI, EVI, and NDWI and reduced the influence of other land features during the downsampling process by sampling 1, 5, and 13 points in each 1 km $\times$ 1 km grid. The importance analysis of this study found that NDVI and EVI were more important than NDWI and far outweighed other meteorological and terrain features. In a study by Michael et al. (2021) in Greece, representing the Mediterranean region, various independent variables including MODIS MOD13Q NDVI products and MODIS MOD09A1 NDWI products were used. The research found that NDWI, as a representation of short-term fuel moisture, is more important than NDVI in fire prediction (Dennison et al., 2005). Bergado et al. (2021), in their wildfire prediction study in Victoria state, Australia, used NDVI, NDII, NDWI, and visible atmospherically resistant index (VARI) calculated from MODIS data as proxies for FMC. The importance ranking conducted in this study found that NDWI and VARI were generally ranked higher than NDVI and NDII.
In addition to optical remote sensing indices, physics-based or data-driven models and microwave data are also can be used in LFMC estimation. For example, Vinodkumar et al. (2021) discovered a lagged relationship between soil moisture variations and corresponding changes in vegetation moisture. By employing soil moisture data, regression analysis, and physical models, they successfully predicted the LFMC for the entire Australia with a lead time of 14 days. Similarly, Lu and Wei (2021) evaluated LFMC (Live Fuel Moisture Content) over the coterminous United States using microwave soil moisture data and employed a time-lagged robust linear regression model. McNorton and Di Giuseppe (2024) further integrated soil moisture data and Leaf Area Index (LAI) to estimate LFMC, and validated the estimates using MODIS data. In addition, Miller et al. (2023) proposed a method using a multi-modal Multi-tempCNN model along with multi-temporal meteorological data, MODIS MCD43A4 data, and climate zone, terrain, and location information. This method enables the estimation of LFMC at a large scale with medium resolution (500m) over a time period exceeding two weeks.
For microwave data, the estimation of LFMC with global coverage in real-time can be achieved by utilizing retrievals of vegetation optical depth through satellite-based passive microwave technology, as demonstrated by Forkel et al. (2023); McNorton and Di Giuseppe (2024) Similarly, (Pelletier et al., 2023) utilized Sentinel-1 backscatter data as proxies of fuel moisture and soil moisture. Rao et al. (2020) designed a physics-assisted neural network to calculate LFMC based on Sentinel-1 microwave backscatter and Landsat-8 optical reflectance data. However, these types of microwave data are susceptible to noise from soil moisture and vegetation types (McNorton and Di Giuseppe, 2024).
2.1.2 Fuel load
Fuel load refers to the total vegetation mass per unit area. Vegetation, especially woody plants, is typically the primary fuel for wildfires, and its condition, quantity, and quality can influence the probability and spreading of fires. Therefore, this parameter can provide a baseline of available biomass in wildfires. Additionally, García-Llamas et al. (2019) found in their study on Mediterranean pine forest ecosystems that fuel load has a greater impact on fire severity compared to fuel moisture content. Thus, estimating fuel load is crucial for fire prediction.
Techniques used to estimate forest structure and biomass include direct and indirect methods of assessment. Direct methods involve destructive sampling of vegetation, which is considered the most accurate form of assessment. Samples are collected on-site, weighed, dried, and then used to estimate biomass volume in the plot. Models have been developed to extrapolate this data over a wider area. However, obtaining a sufficient number of samples can be labor-intensive, costly, and time-consuming.
In contrast to destructive sampling, remote sensing provides an efficient and synoptic means of assessing forest stands. Data captured by satellites over various spatial, spectral, and temporal scales have been used for forest cover and health assessments. Satellites are suitable for covering large areas and conducting multi-temporal analysis. Su et al. (2021), for example, used fractional vegetation cover (FVC) as a proxy for fuel load in their study on wildfire driving factors in parts of southern China’s tropical rainforest. FVC represents the percentage of the total area of the statistical area covered by green vegetation on the ground. In practical applications, MODIS NDVI products are often used as proxies for FVC. Many other research also utilized NDVI from various remote sensing sources as the proxy of fuel load.
However, commonly used sensor bands have broad coverage, including near-infrared and shortwave infrared, which are dominated by vegetation indices associated with water absorption features, masking the absorption features of dry matter such as lignin, starch, and protein. Therefore, Helman et al. (2015) utilized MODIS data and conducted a time series analysis of NDVI data to separate woody and herbaceous plants in Mediterranean forest areas. They derived the long-term average woody NDVI and its trend as proxies for fire fuel.
In addition to NDVI, empirical or physics-based models and data-driven models are also commonly used in fuel load estimation. For instance, Saatchi et al. (2007) used airborne Synthetic Aperture Radar (SAR) images and semiempirical algorithms to estimate the distribution of forest biomass and canopy fuel load in Yellowstone National Park. In addition, Duff et al. (2012) developed a biophysical model that used canopy fuel and NDVI derived from satellite imagery as the main variables to estimate forest fuel quantity. Li et al. (2022b) used the data-driven random forest model to correlate the spatio-temporal features derived from Landsat 8 OLI and Sentinel-1 data with fuel load field measurements. Additionally, the Monte Carlo-driven wildfire simulation model WyoFire (Adhikari et al., 2021) is also an effective method for predicting fuel load and fuel type.
Compared to the fuel load assessment methods based on optical or SAR imagery mentioned earlier, airborne LiDAR can provide more detailed information about vegetation’s three-dimensional structure, such as distinguishing between leaves and branches, and even capturing information about the forest floor. Therefore, it enables more accurate estimation of fuel load at a medium to small scale. For instance, Skowronski et al. (2011) utilized a LiDAR system to estimate crown fuel load in pine-dominated forests of New Jersey, USA, and predicted crown bulk density and crown fuel weight. It is worth noting that ground fuel load is often challenging to capture using optical and microwave signals due to crown obstruction, absorption, reflection, and scattering. As a result, it is often overlooked in most studies, while LiDAR has been proven to have significant advantages in mapping ground fuel. Chen et al. (2017) effectively estimated ground fuel load in Western Australia’s jarrah forests by incorporating LiDAR data with information such as the previous fire year, crown density, elevation, and fuel type. Labenski et al. (2023) further combined Sentinel-2 multispectral data with airborne LiDAR data to extract forest structure and composition information in two forest areas in southwestern Germany. They used a random forest model to predict and map forest surface fuel.
2.1.3 Fuel types
Different fuel types have varying compositions, such as protein, fat, cellulose, and moisture content, as well as different environmental conditions, which influence the likelihood, triggering conditions, and severity of wildfires. In this section, we categorize fuel types into two categories: large-scale and small-scale, to illustrate the influences of different land uses and different vegetation types and classes within the same land use.
Typically, most large-scale research will employ land use maps in wildfire risk prediction. For example, Qadir et al. (2021), in their wildfire risk prediction analysis in Nepal, highlighted vegetation types such as tropical lowland and broadleaved evergreen forests as having the highest fire risk. Jaafari et al. (2019) found in their study on the Zagros ecoregion in western Iran that dry farmlands and forests were the most susceptible areas to wildfires. Instead of forest areas and farmland, Pelletier et al. (2023) studied Canadian boreal peatlands and found that traditional indices designed for upland environments struggle to accurately predict fire risk. Therefore, the study focused more on vegetation moisture content, using features such as Sentinel-1 backscatter data, NDVI, NDII, NDWI, and the Normalized Multi-Band Drought Index (NBDI) to construct long-term (4 weeks) and short-term (6 months) trends and slopes of greenness and wetness as predictors. The study found that the long-term wetness slope of the trend and the short-term NMDI trend had the greatest impact on wildfire prediction using an XGBoost model. Similarly, peatlands that have experienced dry conditions can smolder for extended periods, even under snow cover, becoming a trigger for wildfires.
In forest ecology, fuel load, fuel moisture, and the probability of ignition vary among different tree species. For instance,Bianchi and Defossé (2015) focused on the LFMC (Live Fuel Moisture Content) of different tree species and assessed the negative correlation between LFMC and ignition characteristics under the influence of environmental factors. Furthermore, Gilreath (2006) conducted a wildfire risk study on different tree species and found that lodgepole pine is highly susceptible to wildfires between 10 and 19 years of age. Similarly, Bianchi et al. (2019) compared ignition characteristics between native and exotic conifer species in Northwestern Patagonia, Argentina, and discovered that coniferous forests are more prone to ignition.
2.1.4 Fuel continuity
Furthermore, the role of fuel continuity in wildfire prediction is often overlooked. If combustion generates sufficient heat, adjacent particles such as needles, leaves, grass, and woody debris, or other organic materials, can be heated to their ignition temperature, leading to the sustained spread of fire (Taylor et al., 2013). When fuel continuity is high, the fire can propagate steadily, thereby increasing the risk of wildfires. For example, Qadir et al. (2021), in their study on wildfire risk prediction in Nepal, showed that areas with high forest cover density have a higher risk of fires.Conversely, when canopy discontinuity occurs, the spread of fire can be slowed down or even extinguished (Taneja et al., 2021). Therefore, fuel continuity needs to be considered in research. Various features can be used to represent fuel continuity. For example, Srivastava et al. (2014) suggested that firelines and water bodies, among other isolation measures, reduce the likelihood of wildfire occurrence.
However, due to the limitations of passive optical remote sensing signals and Synthetic Aperture Radar (SAR) in penetrating the canopy, previous research primarily focused on describing canopy characteristics, neglecting the extraction of three-dimensional fuel structure information such as the understory fuel moisture, quantity, and distribution. Although some studies attempted to use multi-angle observations or bidirectional reflectance distribution function (BRDF) inversion to retrieve understory vegetation information, these studies are rarely used in wildfire prediction (Pisek et al., 2015; Canisius and Chen, 2007). Therefore, future research can explore the use of laser radar, such as Single Photon LiDAR, which has better capability to describe understory target features, and investigate how to utilize SAR backscatter data and passive microwave data to extract more precise information on understory target moisture and biomass. Additionally, assessing fuel quantity and types can be relatively challenging in densely forested areas with significant crown overlap. Lastly, compared to other fuel characteristics, the consideration of fuel continuity in most studies is still lacking, especially when using low-resolution remote sensing images or downscaled high-resolution remote sensing images, which may result in the loss of fuel continuity features. Finally, it should be noted that the extraction of fuel characteristics is commonly based on optical images, which are susceptible to interference from clouds, fog, and other noise factors. Currently, there is limited discussion on this aspect.
2.2 Climate and meteorology conditions
Meteorology conditions have been central to nearly every environmental fire danger risk system worldwide. High temperatures, low relative humidity, insufficient precipitation or drought periods, and strong winds have been demonstrated to be critical factors in fire occurrence (Zacharakis and Tsihrintzis, 2023; Keane et al., 2001). On the other hand, climate conditions refer to the long-term average weather conditions and their variations in a region. Climate conditions can control plant distribution, prevailing wind directions, and life cycles (Tavakkoli Piralilou et al., 2022), making them useful for predicting large-scale fire risk over the course of one to several years, even decades.
2.2.1 Meteorology conditions
Firstly, short-term climate conditions primarily focus on research with forecasting periods ranging from quarters, months, weeks, to a few days. They include short-term temperature, wind, relative humidity, soil moisture, precipitation, and lightning, as well as corresponding indices such as the Palmer Drought Severity Index (PDSI) and FWI. Temperature and various humidity conditions can influence fuel condition, such as transpiration and moisture content, which further affect the environmental moisture condition and wildfire risk. Wind direction, wind speed, and other related indices such as mean wind power density (MWPD) and mean wind speed (MWS) can impact fire spread and intensity, making them commonly used parameters in wildfire prediction. Unlike other parameters, lightning can serve as an ignition factor for wildfires.
For example, Tavakkoli Piralilou et al. (2022) used temperature, precipitation, MWPD, and MWS as parameters in their wildfire prediction model for the forest area of Guilan Province, Iran. In another study by Kanwal et al. (2023) conducted in Pakistan, they further utilized monthly average precipitation, transpiration, wind speed, soil temperature, humidity, heat flux, albedo, average surface skin temperature, soil moisture, actual evapotranspiration, water deficit, downward surface shortwave radiation, precipitation accumulation, minimum/maximum temperature, and vapor pressure, totaling 16 meteorology parameters. The assessment of their importance revealed that heat flux, transpiration, and vapor pressure were relatively more important compared to other meteorological parameters, although the differences in importance were not significant. Bergado et al. (2021), in their wildfire prediction study focused on Victoria state, Australia, used seven meteorological factors: daily maximum temperature, daily minimum relative humidity, daily total solar radiation, daily total precipitation, daily mean wind speed, daily mean wind direction, and lightning frequency. The lightning data used was an annual lightning climatology dataset that provided estimates of lightning density for each day of the year. Their research found that total precipitation, lightning density, and surface temperature consistently carried high weights in all models. In contrast,Quan et al. (2023) utilized various meteorological variables, including precipitation, relative humidity, temperature, wind speed, and the Keetch-Byram Drought Index, from the ERA5 dataset, along with other factors, to predict wildfire risk in parts of the western Qinghai-Tibet Plateau in China. The study revealed that relative humidity is the most critical factor associated with forest and grassland fire occurrences in the region. Higher relative humidity leads to higher DFMC, thereby reducing the likelihood of ignition.
2.2.2 Climate conditions
Climate conditions are primarily used to describe annual or multi-year meteorological parameters such as temperature, humidity, precipitation, and evaporation. Similar to meteorology conditions, fuel load, fuel moisture content, and fuel continuity can be considered as functions of climate parameters, making them useful for wildfire prediction (Aldersley et al., 2011; Jaafari et al., 2017). However, the difference between climate and meteorology conditions lies in the fact that climate variables are typically the primary driving factors of large-scale wildfire activity (Jain et al., 2020; Pollina et al., 2013). This is because fuel characteristics respond to climate accumulations before the fire season begins (Chen et al., 2016). Therefore, climate conditions are more suitable for predicting wildfires on large scales and over long periods. Climate conditions can also account for strong seasonal variations in wildfire risk in certain regions (Martell et al., 1989), especially considering the intensifying effects of climate change in recent years, which have raised concerns about their impact on long-term, large-scale wildfire occurrences (Prior and Eriksen, 2013).
For example, Chen et al. (2016) used large-scale sea surface temperature anomalies to predict the severity of the fire season in South America on a seasonal time scale. Their research found that using a single ocean climate index could predict approximately 48% of the global burned area in the three months or longer period before the peak burning month. In a study by Jaafari et al. (2019) conducted in the Zagros ecoregion of Western Iran, they used three climate conditions: annual temperature, rainfall, and wind effect. However, the results showed that only the annual temperature parameter was relatively important. Similarly, Naderpour et al. (2020) studied the Northern Beaches region of New South Wales, Australia, and used three climate factors: annual rainfall, annual humidity, and annual wind speed. Holdrege et al. (2024) conducted research on the big sagebrush region in 13 western states of the United States and used three climate factors: mean annual temperature, annual precipitation, and the proportion of summer (Jun-Aug) precipitation to the annual total precipitation. The study found that areas with low summer precipitation proportions and moderate to high annual precipitation and high temperatures had the highest probability of wildfire occurrence. Similarly, Aldersley et al. (2011) used annual cumulative precipitation, annual lightning frequency, monthly precipitation frequency, monthly average temperature, and climate data from the British Atmospheric Data Centre (BADC) to study the drivers of monthly area burned at global and regional scales. They found that climate factors such as high temperature, moderate precipitation, and dry spells were the most important drivers at the global scale. Mansuy et al. (2019) conducted a study on the United States and Canada and found that climate variables were the dominant controls on area burned for most ecoregions, both within and outside protected areas, surpassing landscape and human factors. The climate data used in their study included mean annual precipitation, mean annual relative humidity, the Hargreave climatic moisture deficit (an annual integrative measure of energy and moisture), and degree-days below 18 °C. Abdollahi and Pradhan (2023) conducted wildfire prediction research in the Gippsland region of Victoria state, Australia, and found that humidity, wind speed, and rainfall were the most influential factors in wildfire prediction. In the study by Zhao and Liu (2021), data from over 600 meteorological stations located in forest, grassland, and shrub-covered areas were used to calculate two drought indices, i.e., Keetch-Bryam Drought Index and Standardized Precipitation Index. These indices were then masked using global-scale data on the PDSI to evaluate long-term wildfire risk in China. The research findings indicated that drought indices showed a stronger correlation with seasonal and annual fire occurrences compared to climate indicators such as temperature and precipitation.
From the above analysis, it is evident that various forms of temperature, wind, humidity, and evaporation data are commonly used as meteorological or climate variables in wildfire prediction models. However, most studies tend to use meteorological variables or climate variables separately, rather than considering them together. Future research should fully consider the impact of meteorological and climate variables at different time scales on the predictive performance of models. In addition, the resolution of meteorological station or reanalysis data used in studies is often relatively low, such as being equal to or greater than 1 km. However, most models operate at a consistent resolution, often downscaling higher-resolution fuel, socio-economic, topographic, and hydrological variables to match the resolution of meteorological or climate data. Therefore, it is worth exploring how to use super-resolution techniques to reconstruct more detailed meteorological or climate data.
2.3 Socio-economic factors
Human activities are important factors in igniting wildfires, with two-thirds of wildfires in the United States being caused by humans Zhai et al. (2003). However, the ignition of fires is stochastic, so many studies introduce socio-economic factors such as distances to infrastructure like roads, power lines, villages, and cities, road density, population density, and land use as proxies for the probability of fire ignition. The basis for this assumption is that wildfire frequency is negatively correlated with road distance and follows a normal distribution with road density (Gilreath, 2006; Martínez et al., 2009). Additionally, urban development can lead to increased fuel load in rural areas, and controlled burning practices in remote areas for agriculture purposes can also increase the likelihood of uncontrolled wildfires (Martínez et al., 2009). Some studies have confirmed these assumptions. For example, Jaafari et al. (2019) found that high wildfire probability is highly associated with road networks in the Zagros ecoregion of Western Iran. Ghorbanzadeh et al. (2019) conducted research in the forestry area of Amol County in the Mazandaran province of northern Iran and similarly found that human activities near major roads had a major influence on wildfire risk. The interaction between human settlements and vulnerable infrastructure areas with high and moderate forest fire susceptibility resulted in high-risk areas in the wildfire risk map.
[…]
Currently, studies on the impact of human activities on wildfire risk primarily focus on incorporating the distribution density of population, roads, and other infrastructure, as well as the straight-line distances from different locations within the study area, as input features for models. However, the consideration of these distances is limited to the simple use of straight-line distance and does not take into account human accessibility. As a result, in areas with rugged terrain, this approach may lead to decreased model performance. Additionally, the current research on human factors mainly focuses on human-induced ignition of wildfires, while rarely considering the inhibitory effects of firefighting facilities such as fire stations, firebreaks, and lookout towers on fire incidents.
2.4 Terrain and hydrological features
Terrain factors used in wildfire prediction studies include elevation, slope, aspect, hill shade, and planar curvature. Terrain indirectly determines mesoscale and microscale climate and airflow (Coen et al., 2013), thereby impacting fuel conditions and wildfire propagation. For example, to some extent, an increase in slope leads to faster fire spread (Ghorbanzadeh and Blaschke, 2018). Steeper slopes result in faster fuel preheating and ignition rates, and compared to flat terrain, the rate of fire spread on a 20° slope increases by about four times (Lecina-Diaz et al., 2014). The slope aspect influences prevailing winds, humidity, solar radiation, and plant species distribution. In many regions of the Northern Hemisphere, north-facing slopes are typically cooler and more moist than south-facing slopes, making south-facing slopes at higher risk of forest fires (Sayad et al., 2019). Curvature and valley depth are geomorphic indicators of soil moisture and vegetation distribution (Kalantar et al., 2020). In larger valleys, downslope-driven cold air pooling has a significant impact on vegetation density and composition (Kiefer and Zhong, 2015). Additionally, elevation can also influence human activities, as lower elevations tend to have higher human population density and are more prone to wildfires compared to higher elevation areas (Guo et al., 2016b). In addition, the main hydrological feature is the distance between fuel and water bodies, soil moisture, and the topographic wetness index (TWI). TWI represents soil saturation and surface runoff rates. Land with higher TWI is more saturated, and high moisture content can prevent fire ignition (Achu et al., 2021).
For example, Ghorbanzadeh et al. (2019) studied the forestry area of Amol County in the Mazandaran province of northern Iran and used slope, slope aspect, and altitude as terrain factors. They found that steep areas were generally more prone to forest fires. Nami et al. (2018) considered terrain factors such as slope, slope aspect, altitude, planar curvature, TWI, and topographic roughness index (TRI) when mapping wildfire risk in the Hyrcanian ecological region in northern Iran. However, their further collinearity analysis revealed significant collinearity between TWI and slope with other parameters, leading to their exclusion from subsequent modeling tasks. A study conducted by Tran et al. (2023) on the island of Oahu in the Hawaiian Islands also utilized terrain and hydrological features such as elevation, slope, aspect, plan curvature, profile curvature, valley depth, TWI, and proximity to rivers. The assessment of their significance indicates that slope, elevation, and TWI are of secondary importance compared to proximity to roads and temperature in relation to fire risk. Gentilucci et al. (2024) also used TWI in a study in the Marche region of central Italy. Their research shows that areas of the terrain with less accumulation of moisture are most susceptible to fires.
Soil moisture information can serve as a complement or substitute for drought indices and is commonly used in wildfire risk prediction models. It can be utilized in modeling fuel moisture and other related factors (Krueger et al., 2022). For example, Bakke et al. (2023) conducted a study in Fennoscandia, Finland, to explore the primary hydro-meteorological factors influencing wildfire occurrence. They found that shallow soil moisture anomalies were the main influencing factor. Similarly, in the study by Vissio et al. (2023) on the prediction of summer burning area in Italy, the ERA5 soil moisture reanalysis data product was used. The research found that using only a single variable, namely surface soil moisture, for predicting wildfire burn area yielded good results. This is because surface soil moisture can serve as a proxy for the combined effects of temperature and precipitation. In the study by Nur et al. (2023) focusing on Sydney, Australia, the MODIS Terra Climate dataset was used, which provided drought indices and soil moisture data. However, due to the low resolution of the dataset (4 km), the importance of soil moisture as a factor was not fully reflected. In research conducted by Yahia et al. (2023), to improve the resolution of soil moisture, temperature vegetation dryness index (TVDI), perpendicular drought index (PDI), optical trapezoid model (OPTRM), and modified normalized difference water index (MNDWI) were used as proxies for soil moisture. These variables, along with a Gaussian Naive Bayes (GNB) classifier, were employed to generate wildfire risk maps. Beyond, the NDVI can be used to estimate soil moisture anomalies (Tucker, 1979; Liu et al., 2010). However, when modeling moisture anomalies using NDVI, it can only directly detect surface soil moisture, and a physical or empirical relationship must be employed to further obtain deep soil moisture. In addition, it is worth noting that the estimation of soil moisture based on optical or SAR remote sensing imagery still faces significant challenges. This is due to the influence of factors such as tree canopy obstruction, absorption, reflection, and scattering, which result in a mixed signal received by the sensor containing information on both soil and fuel moisture (Pelletier et al., 2023).
2.5 Wildfire historical records
Wildfire historical records are essential data in all research studies as they can be used for accuracy validation or model training. The sources and scales of wildfire historical records are diverse. Taking Canada as an example, many provinces, cities, or forest management agencies have their own wildfire historical records. Natural Resources Canada also has its own wildfire historical information data system, such as the Canadian Wildland Fire Information System. In regions where systematic wildfire records are not established, wildfire point identification and historical data recording can be done by combining products from remote sensing satellites such as MODIS, Landsat, and NPP-VIIRS. Specifically, commonly used publicly available global or regional-scale wildfire historical records mainly include the National Burned Area Composite dataset (NBAC), Canadian National Fire database (CNFDB), Instituto Nacional de Pesquisa Espacial database (INPE), Fire Program Analysis Fire-Occurrence Database (FPA-FOD), FireCII, and various MODIS hotspots and thermal anomaly data. It is worth noting that some studies may apply selection criteria to ensure the credibility of these data when using them at a smaller scale.
[…]
MODIS data and products are also commonly used in wildfire risk prediction. For example, Bergado et al. (2021) used the ”Fire History Records of Fires primarily on Public Land” dataset provided by the State of Victoria, Department of Environment, Land, Water & Planning (DELWP) for their wildfire prediction research in Victoria, Australia. They believed that MODIS might overlook fires with smaller burn extents (Hantson et al., 2013). However, Abdollahi and Pradhan (2023) utilized both MODIS MCD64A1 and MOD14/MYD14 fire and thermal anomaly data, along with the NPWS Fire History - Wildfires and Prescribed Burns dataset provided by the NSW Department of Climate Change, Energy, the Environment, and Water, for their wildfire prediction study in the Gippsland region of Victoria, Australia. Ghorbanzadeh et al. (2019) obtained polygonal data for 34 wildfire areas through field surveys and evaluated them using MODIS data for the forestry area of Amol County in the Mazandaran province of northern Iran. Jaafari et al. (2019) and Nami et al. (2018) used historical data, field surveys, and validated them with MODIS hot spot products to enhance the reliability of historical data in their studies on the Zagros ecoregion in western Iran and the Hyrcanian ecological region in northern Iran, respectively. Kanwal et al. (2023) utilized the MODIS Fire Information for Resource Management System (FIRMS) as wildfire historical data for their research in Pakistan and set a confidence threshold of 60% to reduce false alarm errors. Zhang et al. (2021) used the monthly Global Fire Atlas (GFA) product from the Global Fire Emissions Database, covering the period 2003-2016, obtained from the GFA algorithm and 6 MCD64A1 burned area products with a 500 m resolution from MODIS. The GFA has a spatial resolution of 0.25° × 0.25°. They derived four seasonal fire maps, namely March to May (MAM), June to August (JJA), September to November (SON), and December to February (DJF), by averaging the monthly fire products. They analyzed wildfire susceptibility using a binary classification approach and constructed maps indicating the presence or absence of fires by reclassifying the four seasonal fire maps based on the 25th percentile.
3 Data Preparation and Preprocessing
It is worth noting that wildfire prediction often involves the use of multiple types of data, and the values of these data often differ significantly in magnitude, and some data may not be in numerical form. For example, variables like wind direction and land use are typically represented in text form, while fuel characteristics, meteorological variables, terrain, and socioeconomic factors are usually represented numerically. However, the numerical values of these variables can vary greatly, with differences of several orders of magnitude. For instance, the distance to a road and temperature values may differ by several tens or even hundreds of times, and the difference with indices like NDVI could be even larger. Therefore, directly inputting multiple types of data into machine learning or regression models often yields suboptimal results. Additionally, these diverse data types often have different spatial resolutions and temporal scales, which can affect the accuracy and completeness of fuel or environmental information extraction, thereby impacting model performance (Tavakkoli Piralilou et al., 2022). Lastly, different data may have different storage formats, such as meteorological data and historical fire data being stored in tabular or matrix form, fuel data being stored as raster data, and road and human settlement data being stored as vector data in the form of lines or points. However, many studies do not elaborate on the data processing methods used. Given this, this section will outline commonly used data processing approaches, categorizing them into three types: different magnitude data processing, different resolution data processing, and different format data processing.
3.1 Different magnitude data processing
Based on literature research, we found that commonly used methods for processing data with different magnitudes can be categorized into binning, normalization, and trend decomposition. Specifically, Jaafari et al. (2019) categorized each of the 12 variables they used in their study on the Zagros ecoregion in Western Iran into 2-8 classes based on their numerical values. Similarly, Ghorbanzadeh et al. (2019) classified the 16 wildfire-related factors they used in their study on the forestry area of Amol County in the Mazandaran province of northern Iran into 3-9 classes based on expert knowledge and previous references. Salavati et al. (2022) also categorized proxies for terrain, meteorology, and socioeconomic factors into 5-9 classes in their study on Sanandaj city, located in the west of Kurdistan province and the west of Iran. In the modeling process conducted by Bergado et al. (2021), they converted a land cover map into a binary map consisting of 8 features. If a certain land cover category was present, it was assigned a value of 1; otherwise, it was assigned a value of 0. Hong et al. (2019), in their modeling of wildfire susceptibility spatial distribution in Huichang County, China, also transformed continuous landscape indicators into categorical indicators. Rezaie et al. (2023) preprocessed their data by categorizing the 13 predictor variables they used. Liang et al. (2019) recognized that the significant differences in the values of input variables can complicate and prolong the process of finding the optimal solution through gradient descent in deep learning models, thereby affecting the accuracy of predictions. Therefore, they transformed the variables into spatial categories to give them equal weights.
Categorizing variables into classes using specific thresholds requires more expert knowledge and may vary depending on the research context and the types of data used. On the other hand, normalization or trend decomposition, such as Seasonal and Trend decomposition using Loess (STL), can be more reasonable. For example, Malik et al. (2021) normalized precipitation data. Li et al. (2022a) normalized all fire-driving factors to eliminate differences in data ranking and level. Shadrin et al. (2024), when using multimodal data, also normalized the multiple data sources to the range of 0 to 1. In the study by Pelletier et al. (2023) on Canadian boreal forested peatlands, they used STL trend components of the two to four weeks preceding the observation date and the three- and six-month long-term STL trend slopes.
Apart from the significant differences in the magnitudes of different types of data, many studies, especially those employing machine learning and deep learning models, are heavily affected by data imbalance. This is because in the study area, there is a substantial difference in the area between burned and unburned regions. Therefore, many studies address this issue by employing data filtering and augmentation techniques to balance the burned and unburned areas, either in terms of the number of samples or their magnitudes. For example, in the study by Pelletier et al. (2023), when constructing the dataset, they used all wildfire point data along with an equal number of non-wildfire point data, rather than randomly sampling from the entire study area. Similarly, in the study by Kanwal et al. (2023), when training a CNN model, they also selected an equal number of wildfire and non-wildfire point data. Furthermore, Malik et al. (2021) utilized the Synthetic Minority Oversampling Technique to oversample wildfire data, achieving a balanced representation of fire and non-fire data.
3.2 Data with different spatial-temporal resolutions and different formats
In most studies dealing with different temporal and spatial resolutions, a common approach is to resample all the data to the resolution of the lowest-resolution data using popular sampling methods such as nearest neighbor, bilinear interpolation, and cubic spline. Some studies even downsample other driving factors to a single point to match the importance of meteorological station data. However, many studies opt to retain more detailed information and devise alternative sampling methods. For instance, Malik et al. (2021) resampled Landsat 8-derived NDVI, EVI, and NDWI vegetation indices to a resolution of 1 km. To mitigate interference from water bodies and other elements, they used the average value of four corners, four midpoints, centroid, and four diagonal points within the 1 km grid as the sampled data value. Bergado et al. (2021) employed bilinear interpolation to downsample and upsample DEM and LULC data, respectively, to a resolution of 500 m to align with MODIS data. Wang et al. (2021) utilized inverse distance weighting (IDW) to interpolate daily wildfire point data into a monthly gridded burned area of 0.25°×0.25°. Additionally, they employed linear interpolation to interpolate other predictor variables to the same resolution.
In cases where data exhibit temporal sequence gaps, inconsistencies in temporal resolutions among various data sources, or fail to meet specific analysis requirements, the misalignment of different datasets occurs. In such scenarios, interpolation methods are commonly employed to fill in missing data or unify the data at a lower temporal resolution. For instance, Malik addressed missing data at meteorological stations by using the average value of adjacent time series or interpolating with data from the same date in previous years (Malik et al., 2021). Similarly, when establishing models, Bergado et al. (2021) substituted poorly observed solar radiation data with daily averages from previous years at the same time. Pelletier et al. (2023) adopted Kalman interpolation to upsample the temporal resolution of the utilized data to a weekly resolution to meet the requirements of time series analysis techniques.
Different types of data often have significant differences in their storage formats. For example, meteorological station data is typically spatially discontinuous and represented as point data, while satellite remote sensing images are in raster format. Historical wildfire extents, road networks, and river data are commonly represented as vector polygons or polyline data. Therefore, when inputting data into a model, it is necessary to standardize the data with different formats. In many studies, various data from different sources are often unified into a single format, such as a raster format, for model input. For example, Malik et al. (2021) converted vector power line and terrain data into a raster format for processing. Salavati et al. (2022) covered the entire study area by interpolating meteorological station data using inverse distance weighting. Rubí and Gondim (2023), in their study of the Federal District in Brazil, compared interpolation and clustering methods for extrapolating meteorological station data and concluded that interpolation methods better preserved abrupt changes. In contrast, some studies utilize vectors to store various types of data, where a patch of spatial area is represented as a single point, and each element within the vector represents a feature, such as fuel conditions, meteorological conditions, and socioeconomic factors (Gholami et al., 2021).
4 Features Collinearity and Importance Evaluation
The relationship between wildfires and their driving environmental factors has always been a central focus in wildfire prediction research. It is important to consider the relationships between wildfire occurrence and factors such as fuel conditions, climate, meteorological variables, terrain features, and human activities. However, the simultaneous inclusion of highly collinear independent variables in a model can potentially harm its performance (Chen et al., 2019). Moreover, if two or more independent variables are significantly correlated, they provide redundant information in regression analysis, making it difficult to interpret the individual effects of each variable on the dependent variable. Therefore, it is necessary to study the collinearity and importance of different features, as well as the influence of ecoregions and geographical locations on feature collinearity and importance. Commonly used indicators for assessing data collinearity in wildfire risk prediction include Variance Inflation Factors (VIF) and tolerance. In terms of importance evaluation, researchers have utilized the Gini coefficient (Rubí and Gondim, 2023; Nembrini et al., 2018), feature ranking methods (Rubí and Gondim, 2023; Du et al., 2019), SHapley Additive exPlanations (SHAP) method (Abdollahi and Pradhan, 2023; Wang et al., 2021; Lundberg and Lee, 2017), Analytic Hierarchy Process (AHP) (Hong et al., 2019; Saaty, 1977), Information Gain Ratio (IGR) (Rezaie et al., 2023; Quinlan, 1986; Zhang et al., 2019), and ablation study(Shadrin et al., 2024).
4.1 Evaluation of features collinearity
Nami et al. (2018), in their study on wildfire risk mapping in the Hyrcanian ecological region of northern Iran, conducted collinearity analysis using VIF and tolerance. They found significant collinearity between TWI and slope with other parameters, leading to their exclusion from subsequent modeling tasks. Similarly, Jaafari et al. (2019) evaluated the collinearity of various predictors using VIF and tolerance in their study on the Zagros ecoregion in western Iran. They discovered that TWI and slope had VIF values below 5 and tolerance values above 0.2, indicating insignificant collinearity with other predictors (Liao and Valliant, 2012). Hong et al. (2019), while modeling wildfire susceptibility in Huichang County, China, assessed the collinearity of multiple predictors using VIF and tolerance and found no significant collinearity between slope and other features. In a study by Li et al. (2022a) on forest fire driving factors in Yunnan Province, China, it was found that the VIF values for six meteorological factors (daily average air pressure, daily minimum air pressure, daily average temperature, daily minimum temperature, daily average surface temperature, and daily minimum surface temperature) were all above 10, while the tolerance values were all below 0.1, indicating significant correlations. The study also employed Pearson’s correlation algorithm to analyze the correlations of the remaining variables and found a correlation coefficient of 0.77 between surface temperature and air temperature. Similarly, Rezaie et al. (2023), in their wildfire prediction model for Maui, Hawai’i, analyzed the correlations and collinearity among independent variables (elevation, slope, aspect, valley depth, TWI, slope length, plan curvature, distance to the river, distance to the road, NDVI, mean monthly rainfall, mean annual wind speed, and mean annual air temperature) using Pearson’s correlation coefficient, VIF, and tolerance. They found that the strongest correlation was between the average annual temperature and elevation, and the strongest collinearity was between the average annual temperature and distance to the road. However, the collinearity and correlations among all variables were not statistically significant.
4.2 Evaluation of features importance
The feature permutation method, employed in importance evaluation, extracts statistical data on the potential variation of the average absolute error associated with random fluctuations in each feature value. Rubí and Gondim (2023) evaluated the contributions of meteorological, fuel, topographical, and socio-economic factors in models such as ANN, SVM, NB, KNN, LR, LogLR, and AdaBoost, in their study focused on the Federal District, Brazil. They found that analyzing individual variables could lead to contradictory results. For instance, NDVI was more important for models like AdaBoost and ANN, while its impact on other models such as LR and SVM was relatively lower or non-existent. This indicates the necessity of considering multiple variables and the importance of different models in predictive modeling. To address the interpretability challenges of deep learning, Abdollahi and Pradhan (2023) introduced the SHAP method (Lundberg and Lee, 2017) to investigate the contributions of different features to the model. Similarly, Wang et al. (2021) used SHAP to assess the driving factors of wildfires in the contiguous United States, revealing the crucial roles of longitude and latitude in depicting different temporal patterns of fire occurrences. They also discovered the physical relationship between burned area and variables such as vapor pressure deficit, relative humidity (RH), and energy release component (ERC). SHAP has also been applied in wildfire risk mapping studies based on deep learning models (Abdollahi and Pradhan, 2023; Mell et al., 2007) to address the limited interpretability of deep learning. When modeling the spatial distribution of wildfire susceptibility in Huichang County, China, Hong et al. (2019) used the AHP to assign weights to a Logistic Regression model and found that elevation was the most influential indicator for fire ignition, followed by land use, NDVI, and proximity to human settlements. Rezaie et al. (2023), in their wildfire prediction model for Maui, Hawai’i, used the IGR to select the best subset of variables. They found that the distance to roads had the highest IGR value of 0.903, indicating its significant impact on mapping fire-prone areas. The next important variables were average annual temperature (0.724), elevation (0.715), and slope (0.665). Similarly, Zhang et al. (2019), when assessing forest fire risk in Yunnan Province, China, employed IGR to evaluate the importance of different factors. They discovered that temperature and wind speed posed the highest risk, while distance to roads and distance to rivers had the lowest importance. In some cases, ablation experiments can be used to quantitatively assess the importance of different features. For example, Shadrin et al. (2024) evaluated the impact of specific feature values on the final test results by setting them to zero during model testing. Similarly, Chen et al. (2024b) calculated the importance of each feature by measuring the decrease in prediction error when the feature values were shuffled (Breiman, 2001).
5 Prediction Methods
Currently, wildfire prediction methods commonly used include physical models, statistical methods, traditional machine learning methods, and deep learning methods. Statistical methods utilize various regression models to establish the relationships between observed wildfire risk and factors such as fuel conditions, meteorological and climatic conditions, and terrain conditions. By incorporating information related to fuel and environmental variables, statistical methods aim to predict fire risk. These methods can explicitly identify the contributions of different conditions to wildfire risk, and they are relatively mature in terms of modeling techniques and computational requirements. However, statistical methods face limitations due to the strong nonlinearity and randomness of wildfire risk. The predictive capacity of statistical regression models is often limited, as they struggle to capture the complex nonlinear relationships present in wildfire risk. Comparatively, fire danger rating systems employed by many countries for predicting wildfire probability are predominantly based on mathematical modeling principles encompassing fluid dynamics, combustion, and energy transfer and fuel, weather, and topography conditions (Rezaie et al., 2023; Mell et al., 2007). Such models offer better interpretability, generalization capabilities, and are particularly suitable when training data is limited. However, physical models heavily rely on implicit assumptions regarding soil moisture, fuel load, land cover type, landscape structure, air humidity, and related fire behavior, which may adversely affect prediction accuracy (Rezaie et al., 2023; Bar Massada et al., 2011). Additionally, physical models also need to consider explicit modeling of unknown influences such as atmospheric noise and observational instrument effects. On the other hand, statistical methods, such as logistic regression, establish relationships between historical fire data and driving factors, combining this relationship with expert domain knowledge to analyze the contribution of each driving factor to wildfire occurrence risk. These methods are often relatively simple but have limited learning capacity and weak tolerance for handling stochasticity and highly nonlinear wildfire prediction tasks (Li et al., 2022a).
In comparison, traditional machine learning methods, such as RF (Random Forest) and SVM (Support Vector Machines), automatically model patterns and regularities in historical wildfire data through data-driven approaches and manual feature engineering, without explicitly specifying the model type. They can handle nonlinearity with higher interpretability. Compared to traditional machine learning, deep learning models can automatically extract more discriminative independent predictor features, reducing the reliance on manual feature engineering. Moreover, deep learning algorithms have larger parameter capacities, enabling them to handle larger-scale datasets and reducing the risk of model overfitting. Additionally, deep learning models, with their hierarchical structures and nonlinear activation functions, can learn more complex nonlinear relationships and distributions, making them more suitable for complex real-world conditions.
Model evaluation metrics are also a crucial consideration in this study because accurate measurement and assessment of performance are essential when models are applied to real-world scenarios.
5.1 Fire danger rating systems
Due to the strong nonlinearity and randomness of wildfire disasters, it is challenging to establish precise physical models that describe the fire risk related to fuel, weather, terrain, and human activities. In contrast, physical models for wildfire spread prediction are more mature, but they are not the focus of this discussion. Currently, commonly used wildfire danger rating models include the systems proposed by Canada, the United States, Australia, and other countries. Taking Canada as an example, the Canadian FWI system is an integral part of the CFFDRS system. The FWI system is based on the theory of time-lagging and equilibrium moisture content and calculates changes in fuel moisture content based on changes in weather conditions. It then assigns forest potential fire danger levels based on the moisture content of combustible materials at different locations or sizes (Stocks et al., 1989). Additionally, the NFDRS system considers not only fuel and weather but also topography and risk factors in assessing fire risk.
5.2 Statistical models
Statistical models establish relationships between historical wildfire data and driving factors, combining this relationship with expert domain knowledge to analyze the contribution of each driving factor to wildfire occurrence risk. In wildfire prediction, three main types of statistical models have been employed: LR, GLM, GWLR, and other methods. These models aim to capture the dependencies and patterns between wildfire occurrences and various influencing factors. It is worth noting that this section refrains from elaborating on the intricate mathematical definitions of these models, and instead emphasizes their practical application in wildfire risk prediction and performance between different models.
5.2.1 Logistic regressions
Among the various multivariate statistical analysis methods used for predicting natural disasters, logistic regression was once the most popular due to its simplicity in theory, computational efficiency, and its ability to explore the spatial relationships between an event (dependent variable) and a set of independent variables, elucidating patterns of event occurrence and potential influences (Pham et al., 2018). Additionally, the regression coefficients of the logistic regression model can identify the importance of wildfire risk variables, thereby explaining the spatial patterns of wildfire probability in the landscape (Jaafari et al., 2019).
Specifically, Jaafari et al. (2019) conducted a study in the Zagros ecoregion, Western Iran, using two models: the Dempster-Shafer-based evidential belief function (EBF) and LR. They found that LR outperformed EBF significantly. This is attributed to LR’s ability to consider different weights for different features, while EBF assumes equal weights for all variables, which does not align with the actual conditions of wildfire occurrence. Bergado et al. (2021) compared LR with multilayer perception and SegNet (a deep learning network) and found that LR ranked lowest across all three evaluation metrics. Bergonse et al. (2021) utilized logistic regression to predict the burned area during the summer season in Portugal, considering spring meteorological conditions and terrain as independent variables. In a fire risk mapping study in Golestan Province, Iran, Jafari Goldarag et al. (2016) compared LR and ANN, revealing significant differences in accuracy. For instance, in the fire prediction for June 2005, LR achieved an accuracy of 65.76, while ANN achieved 93.49/
With the increase in multi-source data and advancements in computing power, simple LR is no longer sufficient to meet research demands. Several improvements have been made to LR, such as multiple linear regression (MLR) and stepwise linear regression (SLR). The MLR model can consider the influence of multiple independent variables on the dependent variable simultaneously and introduce polynomial terms or other non-linear functional forms to capture non-linear relationships between the variables. SLR, on the other hand, fits the regression model by iteratively adding or removing one covariate at a time based on specified criteria. The aim of this modeling technique is to enhance predictive capability with the minimum number of predictor variables. Compared to LR, both MLR and SLR can accurately capture highly non-linear real-world scenarios. Marcos et al. (2015) developed an MLR model using pre- and current summer drought indices (Standardized Precipitation Index and Standardized Precipitation Evapotranspiration Index) to study the seasonality of summer wildfires in the Mediterranean region (northeastern Spain). The resulting correlation coefficient between drought conditions and summer fire conditions was found to be 0.49. Cai et al. (2019) constructed a tobacco wildfire prediction model in Guiyang County, Guizhou Province, China, using 15 meteorological factors and support vector regression (SVR), with multiple linear regression and stepwise linear regression as comparative models. They found that the SVR model based on support vector machines outperformed LR-based models significantly. Phelps and Woolford (2021) conducted a case study on wildfire incidents in Lac La Biche, Alberta, comparing LR with logistic generalized additive models (LGAM). The results indicated that LGAM, which models highly nonlinear relationships using multivariate smooth functions, achieved better performance.
5.2.2 Generalized linear models
Traditional LR models only consider the linear relationship between independent variables and the outcome, which does not capture the nonlinearity and complexity of real natural disasters. Moreover, multicollinearity among parameters significantly affects the accuracy and reliability of LR models. Therefore, based on the aforementioned comparative studies, it is evident that LR tends to perform poorly. The concept of GLM was developed by Nelder and Wedderburn (1972) and extensively discussed by McCullagh (2019). GLM extends the linear regression framework to handle a broader range of response variables and incorporates non-linear relationships and complex structures. It overcomes the limitations of LR by allowing for different types of response distributions (e.g., binomial, Poisson) and introducing link functions that relate the linear predictor to the response variable. This flexibility enables GLM to capture the inherent nonlinearity and complexity of natural disasters more effectively.
GLM has been widely applied in various fields, including wildfire prediction, to model the relationships between driving factors and fire occurrence. By accounting for different distributions and incorporating appropriate link functions, GLM provides a more comprehensive and accurate framework for understanding and predicting natural disasters, especially in large research areas such as countries, continents, and global scales. For instance, in the global-scale wildfire risk prediction for each month over a year, Song and Wang (2020) compared the performance of neural networks, GLM, and regression tree methods. They found that the neural network approach exhibited superior predictive capabilities. Additionally, Keeping et al. (2024) utilized GLM to predict daily wildfire probability in the contiguous United States. This model demonstrated good performance in predicting the average spatial patterns of fire occurrence, interannual variations in fire counts, and regional changes in wildfire seasonality. The authors believe that although this method was developed for the United States, it is applicable globally.
An important extension of GLM is the Generalized Linear Mixed Model (GLMM) (Breslow and Clayton, 1993), where the response variables are assumed to be mutually independent conditional on a set of latent variables. Ordóñez et al. (2012) used a variation of GLMM called the Generalized Linear Spatial Model (GLSM) to predict lightning-induced wildfires. The results showed that GLSM provided a better fit to the data compared to the standard GLM. The reason behind this improvement is that in non-spatial GLM, the spatial correlation of residuals is not considered, which significantly affects the quality of statistical results.
5.2.3 Geographically weighted logistic regression
Both LR and GLM do not consider the spatial heterogeneity of features. The general assumption of these models is that the relationship between wildfires and their driving factors is constant and stable across space, implying that the relationship between fire ignition and predictor variables is the same throughout the study area (Su et al., 2021). However, this assumption may not hold in reality because predictor variables exhibit high spatial heterogeneity (Li et al., 2022a; Guo et al., 2016a; Chuvieco et al., 2004).
Therefore, the current assumption fails to capture detailed spatial information about the distribution of forest fires and their driving factors, which hinders the detection of fire mechanisms and spatial prediction (Li et al., 2022a). By incorporating spatial components into the model, GWLR accounts for the spatial autocorrelation and enables a more accurate representation of the spatial patterns and dependencies in wildfire occurrence (Chen et al., 2012). This method estimates parameters for each sampling point using weighted least squares, resulting in corresponding estimation coefficients for each sampling point.
Specifically, Li et al. (2022a) employed GWLR in the spatial prediction of forest wildfire driving factors in Yunnan Province, southern China. The evaluation revealed a UC value of 0.902 and a prediction accuracy of 82.9%, indicating good model fit and prediction accuracy. Additionally, the study found that the impact of the same driving factor on forest fires varied across different spatial locations. However, Su et al. (2021) used GWLR in the assessment of wildfire risk in tropical forest systems in southern China and found that its performance was inferior to that of LR, proposing three possible explanations.
5.2.4 Other statistical methods
Furthermore, several other statistical methods, such as the Monte Carlo method, adaptive regression splines, and geostatistical models, have been used either independently or in combination with other models for wildfire risk prediction (Ordóñez et al., 2012; Reyes-Bueno and Loján-Córdova, 2022; Berardi et al., 2023).
5.3 Traditional machine learning models and ensemble models
In the past decade, wildfire modeling has evolved from using statistical algorithms to relatively advanced machine learning methods that can automatically improve through experience (Jain et al., 2020; Oliveira et al., 2021). The main advantages of traditional machine learning algorithms over statistical algorithms are that they do not require prior knowledge of the phenomenon, and they can identify data distributions from large datasets in a non-parametric and nonlinear manner (Oliveira et al., 2021). However, the high dimensionality and complexity of these models make it more challenging to interpret the importance of independent variables compared to traditional statistical regression methods. The use of machine learning methods has become increasingly frequent. Following classification criteria proposed by Jain et al. (2020), commonly used machine learning models for wildfire prediction can be divided into six categories: decision trees, SVM, ANN, Bayesian methods, reward-based methods, clustering methods, and other methods. Some researchers believe that among these models, random forest (a decision tree method) performs better than others in wildfire prediction. For example, Malik et al. (2021) found that in terms of the accuracy of individual models, the random forest algorithm outperformed all other algorithms when experimenting with different target labels and subsets of datasets. However, some researchers argue that different methods have their strengths and weaknesses, and it cannot be concluded that a single model is superior to others (Jaafari et al., 2019; Razavi Termeh et al., 2018). Therefore, some studies combine the predictions from two or even multiple models to improve the final prediction accuracy. Since Jain et al. (2020) provided a detailed review of the applications of machine learning algorithms in wildfire prediction, detection, fuel characterization, and other aspects of wildfire science and management up until 2020, and the subsequent application of traditional machine learning in wildfire prediction is similar to the previous research, this section will not be reiterated.
5.4 Deep learning models
The previously mentioned ML methods are with shallow architectures that do not leverage the spatial patterns inherent in images (Zhang et al., 2018). These classifiers directly classify input data without feature extraction, limiting their ability to extract representative features and improve classification accuracy (Zhang et al., 2019). In recent years, deep learning models have rapidly advanced and outperformed traditional ML models in various tasks, making them important tools for wildfire prediction. However, there is still a lack of comprehensive summaries regarding their application in wildfire prediction, as Jain et al. (2020) only provided a limited overview that included RNN and CNN methods. Therefore, this section systemically reviews the recent progress in deep learning-based wildfire risk prediction models. We categorizes deep learning models used in wildfire risk prediction into RNN methods (including LSTM and GRU) as well as CNN, GCN, GNN, and Transformer, providing a systematic description.
Among these deep learning models, RNN (including LSTM and GRU) is specifically designed to handle time series data, while CNN, GCN, and Transformer models are not inherently designed for sequence data processing. However, many studies based on deep learning or machine learning use grid image data or array data as input and train the models using the wildfire conditions at time t+1 as the ground truth (GT) and the fuel and environmental conditions at time t as independent variables. Therefore, wildfire risk prediction can also be viewed as an image segmentation or classification task, utilizing the aforementioned models to predict the probability of wildfire occurrence at time t+2 based on the environmental and fuel conditions at time t+1.
5.4.1 RNNs
RNN is a neural network structure with recurrent connections used for processing sequential data. It combines the input at the current time step with the output from the previous time step to model and predict sequential data. However, traditional RNNs tend to suffer from the vanishing or exploding gradient problem when dealing with long sequences, making it difficult to capture long-term dependencies. To address this issue, LSTM was introduced. LSTM incorporates three gating units: the input gate, forget gate, and output gate, to control the flow and retention of information. These gating units, with learnable weights, determine whether to pass, forget, or output information based on the current input and the output from the previous time step. This gating mechanism effectively solves the vanishing and exploding gradient problem and allows for better handling of long sequences and capturing long-term dependencies. GRU is similar to LSTM but reduces the number of gating units, simplifying the model structure. Despite its simpler structure, GRU performs comparably to LSTM on many tasks.
For example, Chavalithumrong et al. (2021) used a deep learning network composed of an autoencoder and RNN to predict the burned area of wildfires in parts of North America. The network was trained to predict the observed values and wildfire burned area at time t+1 using the observed values at time t and the state of the RNN. Compared to classical RNN, LSTM is more widely applied in wildfire risk prediction. For instance, Liang et al. (2019) compared the performance of LSTM, RNN, and backpropagation neural network (BPNN) in predicting wildfire size based on meteorological conditions and found that LSTM yielded the best results. Hu et al. (2023) used an autoencoder combined with LSTM to predict the likelihood of wildfires occurring in the United States. Li et al. (2021) introduced an attention mechanism into LSTM to reduce the loss of historical information and improve prediction performance. Similarly, Chen et al. (2024b) proposed a potential wildfire behavior characteristics model by coupling the Bi-directional Long Short-Term Memory and quantile regression methods, which can be used for assessing wildfire behavior characteristics and further enhancing wildfire risk prediction capabilities.
5.4.2 CNNs, GCNs, GNNs, and Transformers
While RNNs have advantages in modeling sequential data by capturing contextual dependencies and temporal correlations, other models such as CNNs, GCNs, and GNNs are also used in wildfire risk prediction tasks due to their stronger local perception capabilities, weight sharing abilities, spatial invariance, and multi-layer feature representation capabilities. In recent years, Vision Transformers, with their strong global feature modeling capabilities, have also been applied to wildfire risk prediction tasks.
For example, Bergado et al. (2021) proposed an AllConvNet, a fully convolutional network, to predict the probability of wildfires burning in the following seven days in Victoria, Canada. Zhang et al. (2019) introduced a CNN combined with fully connected layers for forest wildfire risk modeling in Yunnan, China, based on AlexNet. Their results showed that the proposed method outperformed RF, SVM, MLP, and LR. Huot et al. (2021) treated wildfire risk prediction as an image binary segmentation (semantic segmentation) task and compared Residual Unet and autoencoder, as well as Unet combined with LSTM and autoencoder combined with LSTM. However, they did not find Unet to have significant advantages. Similarly, Shadrin et al. (2024) used a semantic segmentation network called MA-Net to study wildfire spread in the northern regions of Russia and achieved an F1 score of 0.67 in a three-day test.
Ji et al. (2024) utilized the ConvLSTM framework for global wildfire risk mapping, preserving spatial information during the computation to enhance the effectiveness of processing spatiotemporal data. Similarly, Kondylatos et al. (2022) employed the ConvLSTM framework for wildfire risk mapping in the Mediterranean region and found that ConvLSTM outperformed XGBoost and LSTM. Bhowmik et al. (2023) proposed a U-shape LSTM network that achieved an accuracy of over 97% in large-scale wildfire prediction in California, surpassing CNN models.
Some studies address the limitations of CNN in modeling global-scale information and consider the use of GCN and GNN methods in wildfire prediction. Chen et al. (2024a) argued that distant large-scale climate factors and ocean currents can influence wildfire occurrence, which is challenging for CNN to capture. Therefore, they proposed a global wildfire prediction model using GCN in combination with LSTM to improve interpretability. Michail et al. (2024) also emphasized the importance of long time series and large spatial scope information in global wildfire risk prediction. They compared GNN, GRU, and LSTM and found that GNN had stronger consideration of global information, leading to better understanding of complex patterns. The authors highlighted that longer input time series can provide more reliable predictions within different prediction ranges, and integrating spatial information to capture spatiotemporal dynamics of wildfires can improve performance. Therefore, to enhance performance for longer prediction ranges, larger spatial receptive fields need to be considered.
Transformers leverage self-attention mechanisms to capture global dependencies and temporal correlations, enabling them to effectively model complex spatio-temporal relationships in wildfire data. Zerveas et al. (2021) developed a Transformer encoder-decoder for multivariate time series forecasting and demonstrated that the Transformer method outperformed traditional LSTM and GRU models in their task. In the context of wildfire risk prediction, Barmpoutis et al. (2023) used a 360-degree camera combined with a multiscale deformable Transformer to predict the risk of pine forest wildfires in Athens and Thessaloniki, Greece. The results showed that their approach outperformed traditional CNN methods. Furthermore, Miao et al. (2023) proposed an improved Transformer model for forest fire prediction in the Chongli district of Beijing, China. They introduced a windowed attention mechanism to capture and utilize the temporal correlations among multiple variables. Comparisons with LSTM, RNN, and SVM models showed that the improved Transformer achieved better performance. Similarly, Prapas et al. (2023) proposed a teleconnection-driven vision transformer (TeleViT) that integrates fine-grained local-scale inputs with global-scale inputs such as climate indices and coarse-grained global variables for global-scale wildfire area prediction. The experiments demonstrated that the proposed TeleViT method can forecast wildfire areas up to four months in advance and outperforms CNN method and classical vision Transformer model.
6 Future research perspectives
6.1 Data
Regarding the data processing, it is necessary for future research to further enhance the data resolution, consider both short-term and long-term environmental factors, improve the accuracy of fuel condition assessment, enhance the precision of ground truth data, and explore more reasonable methods for constructing training and testing datasets in future studies.
Firstly, the resolution of independent variable features, especially meteorological data, is limited. Many studies downsample all data to match the resolution of meteorological data to maintain spatial consistency, resulting in the loss of detailed information. As mentioned earlier, important parameters such as fuel continuity are significantly affected by spatial resolution. Extrapolating discrete meteorological station data values alone has limited reliability. Similarly, meteorological reanalysis data is also limited by its spatial resolution. Therefore, reasonable methods should be employed, such as super-resolution reconstruction techniques in deep learning, to enhance the spatial resolution of meteorological data and improve the overall spatial resolution of independent variable parameters.
Furthermore, most studies handle fuel and environmental variables by processing data over short or long time scales. When training models, data is often divided into specific time intervals and then randomly shuffled before inputting into the model. Particularly in short-term wildfire probability prediction modeling, researchers often use only a few days, a week, or a few weeks of data to predict the observation period. However, numerous studies indicate that long-term cyclic data, such as wildfire-prone climate data, can also provide support for wildfire ignition research.
Thirdly, the assessment of fuel, especially fuel moisture content and fuel load, is predominantly carried out in two dimensions. In areas with tree cover, especially with high forest density and complex structure, the information on understory fuel, surface fuel load, and humidity is often overlooked because visible and infrared signals, as well as radar signals like SAR, have difficulty penetrating the canopy. Furthermore, the estimation accuracy of fuel attributes (such as canopy base height, depth of fuel bed, and dead fuel loads) is greatly influenced by the lack of field surveys. Therefore, future research should consider the use of three-dimensional laser radar data that provides better penetrations of the canopy, such as single-photon laser radar, to estimate understory fuel conditions and the forest environment. Furthermore, when studying short-term wildfire risk in time series, the assessment of fuel conditions based on two-dimensional passive optical and passive microwave remote sensing proxies can be susceptible to cloud, fog, and smoke occlusions, potentially resulting in signal loss. It is important to further investigate methods for establishing consistency between active remote sensing signals or physical models and other passive signals to infer temporally and spatially continuous fuel conditions.
Following, all supervised statistical regression, traditional machine learning, and deep learning methods rely on historical wildfire point data as ground truth for training models. However, the existing publicly available historical wildfire data has several limitations, such as the neglect of small fire points and low confidence levels. As a result, many studies have resorted to simply applying credibility thresholds and land use filtering to satellite thermal anomaly data in order to select historical burn areas. However, this filtering approach may lead to the overlooking of many ignition points, potentially resulting in systematic underestimation by subsequent trained models. Therefore, future research should prioritize the screening of low-confidence and small-scale wildfires to ensure data quality. Furthermore, the dataset should focus on tracking the dynamic changes in wildfire development to determine the temporal variations of ignition points and wildfire progression, thus providing a solid data foundation for broader applications. Hence, further research is required to explore how to effectively utilize ground inventory historical wildfire data and multi-source active and passive remote sensing data to establish high-quality, high spatial resolution, long-term, and large-scale historical wildfire datasets.
Finally, in most studies, when constructing training and testing sample sets, the areas where wildfires have occurred are labeled as 1, while the areas where wildfires have not occurred are labeled as 0. This is achieved by random sampling within the study area to ensure that the ratio of wildfire samples to non-wildfire samples is close to or equal to 1:1. However, this approach leads to two issues. Firstly, historical wildfire data is not always accurate. Therefore, using the probabilities of 0 and 1 assigned based on historical data does not accurately represent the true occurrence of wildfires. For example, some studies only consider wildfires with MODIS confidence values above a certain threshold, while other publicly available data products may overlook small-scale wildfires. Therefore, the absence of wildfires in an area does not necessarily mean that the probability of wildfire occurrence is zero, making it unreasonable to assign a probability of 0 to such areas. Secondly, most data-driven algorithms set the ratio of wildfire samples to non-wildfire samples as 1:1 to address the impact of class imbalance on algorithm performance. However, this ratio does not align with the actual distribution characteristics of wildfires. Currently, there has been limited attention given to the issue of data imbalance in wildfire risk prediction research. There is a lack of systematic studies on the sampling range of non-wildfire areas, the influence of different confidence levels in historical records on sample ratio settings, and the effectiveness of oversampling and undersampling methods in wildfire risk prediction.
6.2 Methods
Currently, several tasks need to be undertaken to enhance the effectiveness of wildfire risk prediction models. These tasks include improving model performance, incorporating the element of randomness in wildfire ignition, integrating relevant knowledge into the development of prediction models, accurately predicting real wildfire risk probabilities, evaluating the confidence and uncertainty associated with model predictions.
Firstly, machine learning remains the main approach for wildfire prediction research in recent years. The main reasons for this are that machine learning models have stronger interpretability compared to deep learning and do not require extensive data for training. Additionally, machine learning performs better than traditional statistical regression models. However, machine learning models have limited parameter capacity compared to deep learning and are not suitable for large-scale data input, feature extraction, and modeling of high-resolution, large-scale, and long-time series data with strong randomness and nonlinearity. Additionally, it is evident from data analysis that both historical fire point data and remote sensing data capable of inverting fuel characteristics and environmental conditions are abundant, providing a solid foundation for training the deep learning model. This kind of insufficient utilization of data has resulted in a bottleneck in the predictive performance of the current wildfire prediction model. Therefore, it is important to explore how to use state-of-the-art time series forecasting or image processing deep learning algorithms to automatically extract multiple deep features and improve the accuracy of wildfire prediction models. Hence, an important topic is how to fully utilize various data sources and design a reasonable foundation model that can be applied to wildfire prediction at different spatial and temporal scales globally.
Secondly, wildfire ignition is highly random. Typically, research attributes it to socio-economic factors such as agricultural burning, abandoned ignition sources, as well as natural factors like lightning. Socio-economic factors are often represented by population density, infrastructure density, and distance to infrastructure. However, the relationship between these factors and wildfire ignition is not very strong, as it does not take into account factors such as human accessibility to fuels and the randomness of human behavior. It is important to note that wildfire ignition is the end result, and researchers should focus more on fuel and environmental variables, allowing the model to produce more false-positive predictions to prevent wildfire ignition. Developing a practical and reasonable evaluation system, training and evaluating models that are more tolerant of false-positive predictions, should also be considered in future research.
Thirdly, the commonly used methods currently employ various remote sensing data as proxies for fuel, environment, and triggering conditions. However, wildfire behavior is a complex physical process that involves fuel burning, heat transfer, and fluid dynamics, among others. It is challenging to effectively model the physical behavior patterns of wildfires, such as the potential rate of spread (ROS), solely using remote sensing data and deep learning methods. Therefore, it is crucial to couple wildfire combustion physics models, proxy data, and deep learning models to gain a comprehensive understanding of wildfire behavior. In addition, introducing knowledge into deep learning models, rather than solely relying on data-driven approaches, can effectively reduce the reliance on training data and enhance the interpretability of the models. In turn, it allows for targeted interventions in fuel or environmental management, thereby reducing the risk of wildfires.
Furthermore, it is important to note that the predicted results of most models mentioned above are not true probabilities since they need to account for data imbalance, which requires resampling or augmentation techniques to alter the true data distribution. Additionally, the predicted probabilities of some machine learning models may not necessarily equal the true probabilities. For instance, in XGBoost, the predicted class probabilities are obtained through an inverse logarithmic transformation of the sum of leaf weights at the terminal nodes. Similarly, the output of an SVM model is determined by the position of a data point relative to the decision boundary rather than directly representing a probability. Likewise, the probability estimates of deep learning models may not be perfectly calibrated. Many studies simply employ the sigmoid function to map the prediction results to the range of (0,1), without explicitly modeling the correspondence between predicted probabilities and the true likelihood of a sample belonging to a particular class. Addressing the issue of calibration and obtaining well-calibrated probabilities from machine learning and deep learning models is an active area of research. Various calibration techniques, such as Platt scaling, isotonic regression, and temperature scaling, have been proposed to improve the calibration of predicted probabilities. However, these methods have been scarcely explored in the context of wildfire prediction tasks.
In addition, Confidence and uncertainty evaluation of model predictions: It is crucial to evaluate the confidence and uncertainty of wildfire risk predictions. Users need to understand the reliability and limitations of the model outputs. However, existing models often lack robust methods for quantifying and communicating prediction uncertainties. Incorporating techniques such as probabilistic modeling, ensemble methods, or Bayesian approaches can enhance the reliability and interpretability of predictions by providing measures of confidence and uncertainty.
7 Conclusion
In order to perform wildfire prediction, researchers have utilized various data sources and proposed different models and methods. However, there is still a need for a more comprehensive review of these studies. Therefore, this paper has provided a comprehensive review of the data and methods, aiming to discuss the classification and frequency of various data sources and methods, provide support for selecting appropriate data and processing workflows for different ecological areas, and summarize the effectiveness of different models and commonly used model evaluation methods under different study conditions. Additionally, the paper also explores the issue of selecting appropriate time scales for specific scenarios.
Firstly, the establishment of wildfire prediction models requires multiple input data sources and ground truth for training and validation. The input data includes fuel, meteorological and climatic conditions, terrain variables, hydrological characteristics, and socio-economic factors, among others. Researchers choose different data sources and design corresponding data preprocessing and processing workflows based on the characteristics of the study area and data availability. It was found that most studies prioritize meteorological and climatic parameters, followed by fuel and terrain conditions, and finally, socio-economic and hydrological factors. Various methods are used to evaluate the collinearity and importance among different features, but the collinearity and importance between different data sources vary significantly depending on the study area. Ground truth data, including historical wildfire records from field measurements and remote sensing imagery, among others, form the basis for model training and evaluation. This paper collected commonly used open-source data features as well as data post-processing and filtering methods.
Another key aspect of wildfire prediction models is the selection of algorithms and the design of accuracy evaluation systems. Currently, commonly used models include physical models, statistical models, traditional machine learning models, and deep learning models. Physical and statistical models have been widely applied in wildfire prediction, while the application of deep learning models has gained attention in recent years. This paper classified deep learning models, including RNN, CNN, GCN, GNN, and Transformer, among others. The comparison shows that deep learning methods significantly outperform traditional machine learning and statistical models.
Finally, this paper summarizes and looks ahead to the current research gaps in terms of data and methods. This includes improving data resolution, paying more attention to large-scale spatial and temporal data, utilizing multi-source data to enhance fuel and environment extraction capabilities, and establishing more reliable historical wildfire record datasets. In terms of methods, it is important to combine the latest deep learning algorithms with physical models to overcome the accuracy bottleneck in wildfire prediction.

### v2

@@
 Wildfire Risk Prediction: A Review
-Abstract
+Wildfire Risk Prediction: A Review
+Zhengsen Xu Jonathan Li Sibo Cheng Linlin Xu
+Résumé
 Wildfires have significant impacts on global vegetation, wildlife, and humans.
@@
 Currently, there is limited discussion on this aspect.
-2.2 Climate and meteorology conditions
+2.2 Climate and weahter conditions
 Meteorology conditions have been central to nearly every environmental fire danger risk system worldwide.
@@
 Climate conditions can control plant distribution, prevailing wind directions, and life cycles (Tavakkoli Piralilou et al., 2022), making them useful for predicting large-scale fire risk over the course of one to several years, even decades.
-2.2.1 Meteorology conditions
+2.2.1 Weather conditions
 Firstly, short-term climate conditions primarily focus on research with forecasting periods ranging from quarters, months, weeks, to a few days.
@@
 For example, Bergado et al.
-(2021) used the ”Fire History Records of Fires primarily on Public Land” dataset provided by the State of Victoria, Department of Environment, Land, Water & Planning (DELWP) for their wildfire prediction research in Victoria, Australia.
+(2021) used the "Fire History Records of Fires primarily on Public Land" dataset provided by the State of Victoria, Department of Environment, Land, Water & Planning (DELWP) for their wildfire prediction research in Victoria, Australia.
 They believed that MODIS might overlook fires with smaller burn extents (Hantson et al., 2013).
@@
 Model evaluation metrics are also a crucial consideration in this study because accurate measurement and assessment of performance are essential when models are applied to real-world scenarios.
+5.1 Quantitative Statistics of Reviewed Papers
+To provide a quantitative overview of commonly used methods, we analyzed 243 papers to determine the frequency of methods and features used.
+Additionally, we examined the spatial distributions of this research to highlight the fairness in research coverage.
+5.1.1 Frequency of Different Methods
+In this analysis, we categorized 247 models into statistical-based (S-based), traditional machine learning-based (TML-based), deep learning-based (DL-based), ensemble (E-based), experience-informed (E-informed), physics-informed (P-informed), and other methods, as illustrated in Fig.
+1.
+Overall, the most frequently used method is traditional machine learning, appearing 115 times in our analysis.
+This is followed by deep learning methods and statistical-based methods, which were used 41 and 38 times, respectively.
+Figure 1: The frequency of different methods in wildfire risk and spread prediction.
+Within the traditional machine learning-based methods, random forest is the most preferred method, used 31 times in our analysis.
+This preference may stem from two aspects.
+Firstly, 31.25% of comparative studies indicate that random forest achieved state-of-the-art (SOTA) performance compared to other methods.
+Additionally, random forest offers explainability, which helps in identifying the importance of different features and aids decision-makers in understanding and intervening in factors influencing wildfire risk.
+Notably, most comparative studies did not include deep learning methods.
+Other traditional machine learning methods such as boosting, MLP, and SVM-based methods achieved SOTA performance 5, 2, and 1 times, respectively.
+Other methods, including decision tree, ANN, and Bayesian methods, were used approximately 11 times.
+Statistical-based methods are also widely used due to their simplicity and ease of explanation.
+Among these methods, logistic regression is the most popular, utilized 20 times.
+GLM and WOE were each used 4 times.
+Additionally, 10 other regression-based methods were explored in various wildfire prediction studies.
+Compared to shallow models, deep learning methods, particularly CNN and RNN, are also commonly used due to their high accuracy resulting from automatic deep feature extraction, despite their disadvantage in explainability.
+In our analysis, these two models were used 17 and 14 times, respectively, especially in combinations such as CNN and LSTM.
+Graph-based and vision transformer-based methods were used 4 and 1 times, respectively, to extract long-range dependencies given the complex fuel and environment of wildfire-prone areas.
+It is noteworthy that autoencoder-based methods were used 3 times, likely due to the lack of large-scale annotated wildfire data for deep learning model training.
+Another approach to predicting wildfires is to combine various models and knowledge to enhance simulation, improve model performance, or reduce the data dependency of machine learning methods.
+In this section, we divided the remaining methods into four groups: combinations of different models, knowledge-informed methods, physics-informed models, and other models.
+Ensemble methods are commonly used to predict wildfires.
+This research often involves combining several machine learning methods either to achieve final predictions or to form a complex wildfire risk prediction system.
+One common approach is to weight the predictions from different independent models to enhance accuracy (Ghorbanzadeh and Blaschke, 2018).
+For example, Rezaie et al.
+(2023) combined the group method of data handling (GMDH) and convolutional neural network (CNN) algorithms with biogeography-based optimization (BBO) and ant colony optimization (ACO) algorithms to improve predictive performance.
+Another approach involves building complex wildfire risk prediction systems.
+For instance, Ghorbanzadeh et al.
+(2019) combined machine learning (ML) methods with geographic information system multi-criteria decision-making (GIS-MCDM) methods to establish a forest fire susceptibility index and a social/infrastructure vulnerability index.
+Compared to fully data-driven machine learning methods, 18 methods incorporated knowledge, such as observation bias, inductive bias, and learning bias (Karniadakis et al., 2021), to enhance model performance or reduce the reliance on large training data sets.
+For example, Troumbis et al.
+(2023) combined the theory of wildfire self-organized criticality (SOC) with modern statistics to provide a realistic explanation of the phenomenon’s evolution over recent decades and predict trends in Greece.
+Similarly, Bayani et al.
+(2023) used three-dimensional nonlinear vibration equations of power lines to generate a data set considering physical, structural, and meteorological parameters to address the problem of wildfires caused by power lines in California, USA.
+Machine learning techniques were then used to develop an agent model to quantify the risk of a single power line igniting a wildfire under extreme weather conditions.
+Additionally, Prestemon et al.
+(2002) used a fire risk function rather than a data-driven model to explore the relationship between wildfires and vegetation management, human land use, and climate factors in Florida counties.
+Physics-informed models are commonly used in wildfire risk prediction tasks.
+Unlike knowledge-informed methods, these models are based on physical estimations rather than incorporating physical knowledge into other basic models.
+These methods are often used to directly estimate wildfire risk or simulate wildfire spread.
+For example, Liu et al.
+(2010) used a regional climate model to predict conditions in the northwestern United States.
+This model estimated precipitation, humidity, temperature, wind, soil, and vegetation conditions, which were then used to evaluate wildfire potential.
+Similarly, Mitsopoulos et al.
+used the minimum travel time algorithm to simulate wildfire risks under three different weather and fuel moisture conditions to better understand the interactions between wildfires and environmental and fuel conditions in Greece.
+Additionally, Gonzalez-Olabarria et al.
+(2023) employed an improved version of the Cell2Fire simulator, Cell2Fire_SB, to simulate the generation and spread of crown fires based on fuel models from Europe and the United States.
+Moreover, Lattimer et al.
+(2020) introduced machine learning into computational fluid dynamics (CFD)-based wildfire simulation, showing that prediction efficiency was 2-3 orders of magnitude faster than CFD simulations.
+5.1.2 Research area distribution
+Fig.
+2 illustrates the geographic distribution and quantity of wildfire prediction research globally.
+The figure indicates that North America and Asia are the most concentrated regions for wildfire prediction studies.
+The United States has the highest number of studies, totaling 71, followed by China and Australia, each with 18 studies.
+Additionally, Canada, Iran, and India exhibit considerable research activity, with 11, 14, and 6 studies, respectively.
+Figure 2: The spatial distribution of included research.
+In Europe, the research is primarily concentrated in Western and Southern Europe.
+Spain, France, Greece, and Portugal have a notable number of studies, with 8, 5, 7, and 9 studies, respectively.
+Other countries such as Italy, Sweden, Serbia, and Turkey also have a smaller number of studies.
+Research on the African continent is relatively sparse, with South Africa, Algeria, Morocco, and Senegal each having one study.
+Similarly, South America has a limited number of studies, with Brazil and Chile each contributing one study.
+In Oceania, aside from Australia, no other countries have been recorded as having studies.
+Furthermore, there are several studies categorized as global, totaling 18, and studies focused on the Mediterranean basin, totaling 6.
+These global and regional studies highlight the complexity and diversity of wildfire prediction across different climate zones and ecosystems.
+Given the geographic distribution and quantity of research, future studies could focus on expanding research regions, promoting cross-regional studies, and enhancing data sharing and model validation.
+Specifically, although North America and Asia have concentrated research efforts, other regions such as Africa, South America, and parts of Europe have significantly fewer studies.
+Future research should aim to strengthen wildfire prediction studies in these underrepresented regions to address the geographic distribution gaps.
+Additionally, global and Mediterranean basin studies suggest that cross-regional research is crucial for understanding wildfire behavior under different ecosystems and climatic conditions.
+Enhancing international collaboration to conduct more comprehensive cross-regional studies can provide more robust prediction models and methods.
+Finally, many current studies rely on region-specific datasets, limiting the generalizability and applicability of the models.
+Future efforts should focus on promoting data sharing and openness, establishing unified data platforms, and conducting model validation across different regions to improve the reliability and applicability of prediction models.
 5.2 Fire danger rating systems
@@
 In terms of methods, it is important to combine the latest deep learning algorithms with physical models to overcome the accuracy bottleneck in wildfire prediction.
+Références